# Review and Prospect of Algebraic Research in Equivalent Framework between Statistical Mechanics and Machine Learning Theory

### v2

@@ 4 Conclusion
In this paper, for theoretical physicists who are interested in artificial intelligence, we introduced mathematical equivalence between statistical mechanics of random Hamiltonian and machine learning theory. Professor Huzihiro Araki indicated that algebraic concepts such as operator algebras play important role in statistical mechanics and quantum field theory. In this paper, we proposed that, also in machine learning theory, algebraic research is useful in both fundamental and practical fields. For example, such researches are becoming more important in developing statistics and artificial intelligence.
@@ Acknowledgment
The author would like to express his sincere gratitude to Professor Huzihiro Araki who is a pioneer founder of algebraic research in both statistical mechanics and quantum field theory.

### v3

@@ 4 Conclusion
In this paper, for theoretical physicists who are interested in artificial intelligence, we introduced mathematical equivalence between statistical mechanics of random Hamiltonian and machine learning theory. Professor Huzihiro Araki indicated that algebraic concepts such as operator algebras play important roles in statistical mechanics and quantum field theory. In this paper, we proposed that, also in machine learning theory, algebraic research is useful in both fundamental and practical fields. For example, such researches are becoming more important in developing statistics and artificial intelligence.
@@ Acknowledgment
The author would like to express his sincere gratitude to Professor Huzihiro Araki who is a pioneering founder of algebraic research in both statistical mechanics and quantum field theory.